In a bottle of soda, there is 0.29 mole CO2. How many grams of CO2 are in the bottle?
1
Expert's answer
2020-10-12T13:33:40-0400

n = m/M

m = n * M;

M(CO2) = 12 + 16*2 = 44 g/mol

m = 0,29 mol * 44 g/mol = 12.76 g